Using the mathematical model: to calculate velocity of a car accelerating from rest in straight line. Equation is v(t) = A (1-e ^ -t/t max speed) the ^ means to the power of. v(t) is the instantaneous velocity of the car (m/s), t is the time in seconds, tmaxspeed is the time to reach the maximum speed in seconds and A is a constant. We are given the information that the car has a t(0-28m/s) of 2.6s, a t(400m) of 10.46s, and a tmaxspeed of 7s (S stands for seconds in this instance) 1. Identify the units of the coefficient A, 2. physical meaning of A, 3. velocity of the car at t=0 and 4. asymptote of this function as t ⮕ ∞?
1
Expert's answer
2020-01-07T10:24:33-0500

"v(t)=A(1-e^{-t\/t_o})"

"A=constant"

"t=time" (in seconds)

"t_o=" time to reach the maximum speed

"t_o=7s"

1.The coefficient A has a unit same as the velocity, that is, m/s.

2.Terminal Velocity(as t tends to infinity, velocity tends to A)

3."v(t)=A(1-e^{-t\/t_o})"

"A(1-e^{-(2.6\/7)})=-28"

"A(0.31)=-28"

"A=-28\/0.31=-90.322m\/s"

"v(0)=A(1-1)=0m\/s"

4.As t tends to infinity,v(t) tends to the terminal velocity equal to A.

The asymptote of this function is "y=-90.322" .